OpenStudy (anonymous):

Who loves word problems! plz help me. Jamie spent twice as much money as Tina in a clothing store. Ariel spent half as much as Tina did. Christina spent half as much money as Ariel did. Ella spent 12 dollars, which was half as much as what Christina spent. How much did each girl spend?

OpenStudy (mertsj):

x = amount tina spent 2x= amount jamie spent 1/2 x = amount ariel spent 1/2(1/2 x)= amount Christina spent 1/2 (1/2)(1/2 x) = amount Ella spent = 12

OpenStudy (anonymous):

Define some variables. J = money spent by Jamie T = money spent by Tina A = money spent by Ariel C = money spent by Christina E = money spent by Ella Jamie spent twice as much money as Tina in a clothing store. Ariel spent half as much as Tina did. Christina spent half as much money as Ariel did. Ella spent 12 dollars, which was half as much as what Christina spent. J = 2T A = 1/2 *T C = 1/2 *A E = 12 = 1/2 * C

OpenStudy (mertsj):

So: \[\frac{1}{8}x=12\] \[x=96\]

OpenStudy (mertsj):

You can figure out the rest.

OpenStudy (anonymous):

E = 12 1/2C = 12 -> C = 24 24 = 1/2 A -> A = 48 48 = 1/2 T -> T= 96 J = 96*2 -> J = 48

OpenStudy (anonymous):

thanks !